# PCL-Indexability and Whittle Index for Restless Bandits with General Observation Models

### v3

@@ 2.6 Optimality of Index Policy
In fact, by Theorem 1, if the initial state is $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, then for any $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, $b(\{0\}\cup\Omega)=\frac{1}{1-\beta}-T_{j}^{\Omega^{c}}$. As for $b(\Omega)=\inf\limits_{\pi}\{\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\}$, when we always active the auxiliary arm, then $\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}=0$. Combining this fact with $b(\Omega)\geq 0$, we have $b(\Omega)=0$. Next, we only need to consider whether state 00 is in $\Omega_{*}$ or not:
@@
(ii) If $0\notin\Omega_{*}$, then $\exists N$, when $n>N$, $0\notin\Omega_{2n-1}$. In this case, $b(\Omega_{*})=b(\Omega_{2n-1})=0$. In summary, we obtain $b(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n+1})$. Similar arguments show that $\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n})\leq b(\Omega_{*}^{-})$. $\Box$
@@
From the weak duality theorem and above discussion, we know that the optimal policy is the priority policy given by $\{\lambda_{\boldsymbol{\omega_{i}}}\}$. It follows from $A_{0}^{\{0\}\cup\Omega}=1$ and $W_{0}^{\{0\}\cup\Omega}=\lambda$ that the priority index for state 0 is $\lambda$. Therefore, for state $\boldsymbol{\omega_{i}}$, when the system’s subsidy $\lambda=\lambda_{\boldsymbol{\omega_{i}}}$, state $\boldsymbol{\omega}$ is precisely at the critical point for taking active or passive action under the optimal policy. Therefore, this two-arm problem is Whittle indexable, and the Whittle index for state $\boldsymbol{\omega_{i}}$ is $\lambda_{\boldsymbol{\omega_{i}}}$. $\Box$
@@ 3.1 $T$-Step Approximate State Space
If the repeated states are also taken into account, then $|\Omega(T|\boldsymbol{\omega})|=\frac{(L+1)^{T+1}-1}{L}$. In this case, the number of states will increase exponentially with the number of time steps $T$. However, in practical situations, the phenomenon of state repetitions often occurs due to the long time horizon of the underlying stochastic processes in consideration. In other word, along the process of belief state updates, the updated state may have already appeared within a small number of steps from the start or have a short Euclidean distance to a previous state. In such cases, we can assume that the states encountered within a finite number of steps approximate the true belief state space reasonably well. In this sense, the cardinality of the $T$-step state space can be significantly reduced. Formally, we recursively define the $T$-step approximate state space as follows.
@@
For any $\varepsilon>0$, belief state $\boldsymbol{\omega}$ and belief state space $\Omega$, we say $d(\boldsymbol{\omega},\Omega)\geq\varepsilon$, if $\forall\boldsymbol{\omega}^{\prime}\in\Omega$, we have $||\boldsymbol{\omega}-\boldsymbol{\omega}^{\prime}||\geq\varepsilon$. Define
$\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})=\begin{cases}\{\mathcal{B}_{l_{1}}\cdots\mathcal{B}_{l_{T}}\boldsymbol{\omega}:0\leq l_{1},\cdots,l_{T}\leq L\text{ and}\\ d(\mathcal{B}_{l_{1}}\cdots\mathcal{B}_{l_{T}}\boldsymbol{\omega},\tilde{\Omega}(T-1,\varepsilon|\boldsymbol{\omega}))\geq\varepsilon\}\cup\tilde{\Omega}(T-1,\varepsilon|\boldsymbol{\omega}),&T\geq 2\\ \{\mathcal{B}_{l}\boldsymbol{\omega}:0\leq l\leq L\text{ and }d(\mathcal{B}_{l}\boldsymbol{\omega},\{\boldsymbol{\omega}\})\geq\varepsilon\}\cup\{\boldsymbol{\omega}\},&T=1\end{cases}.$
@@
Under the definition of $T$-step approximate state space, numerical calculations show that $|\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})|$ is likely to be much smaller than $|\Omega(T|\boldsymbol{\omega})|$, which will greatly reduce the actual number of states that need to be considered for the problem. The following example illustrates the above concept.
Example 1.
Set $\textbf{P}=\begin{pmatrix}0.8&0.2\\ 0.2&0.8\end{pmatrix},\textbf{E}=\begin{pmatrix}0.8&0.2\\ 0.2&0.8\end{pmatrix},\textbf{R}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$ and $\boldsymbol{\omega}_{0}=(0.6,0.4)$. Assume the upper limit of norm error $\varepsilon$ is $10^{-3}$. After belief updates based on observations, then the number of belief states $|\tilde{\Omega}(6,10^{-3}|\boldsymbol{\omega})|$ in the 6-step approximate space space and the number of states $|\Omega(6|\boldsymbol{\omega})|$ in the 6-step state space are 309 and 1093. If the case under consideration is $\varepsilon=1e-4$ and $T=10$, we have $|\tilde{\Omega}(10,10^{-4}|\boldsymbol{\omega})|=3550,|\Omega(10|\boldsymbol{\omega})|=88573$. These two examples show that in the approximate sense, the number of states to be considered under the update rule is far less than the number of states in the precise sense, which practically allows us to transform the infinite-state space problem into a finite-state space problem. Furthermore, if the obtained reward is the only observation and the upper limit of norm error $\varepsilon$ is $10^{-3}$, we have $|\tilde{\Omega}(6,10^{-3}|\boldsymbol{\omega})|=97$ and $|\tilde{\Omega}(10,10^{-3}|\boldsymbol{\omega})|=158$.
Fact 1.
@@
For any given $\varepsilon>0$, $\exists T$, such that $\forall t>T$ and $\boldsymbol{\omega}_{t}\in\Omega(t|\boldsymbol{\omega})$, $\exists\boldsymbol{\omega}_{T}\in\tilde{\Omega}(T|\boldsymbol{\omega})$, such that $||\boldsymbol{\omega}_{t}-\boldsymbol{\omega}_{T}||<\varepsilon$, where |||||$\cdot$||| | ⋅ | | is the Euclidean norm.
Fact 2.
@@
$\forall\varepsilon>0$, $\lim\limits_{T\rightarrow+\infty}|\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})|<+\infty$.
Facts 1 and 2 can be easily proved by the compactness of the belief space $\Omega$ (a compact simplex). By these facts, for a sufficiently large number $T$ of time slots, the belief states can be well approximated by a finite set of belief states within $\tilde{\Omega}(T|\boldsymbol{\omega})$.
@@ 3.2 AG Algorithm for Approximate State Space
Now we are ready to incorporate the approximate state space $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$ into the AG algorithms introduced in Niño-Mora (2002) based on Theorem 12 and Facts 1 and 2. Specifically, for the state update, if the resulting belief state $\boldsymbol{\omega}$ does not fall within $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$, then we look for the belief state $\boldsymbol{\omega^{\prime}}$ in $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega_{0}})$ that has the closest Euclidean distance to $\boldsymbol{\omega}$, and use the Whittle index $\lambda_{\boldsymbol{\omega^{\prime}}}$ of $\boldsymbol{\omega^{\prime}}$ as the Whittle index of $\boldsymbol{\omega}$ or adopt a linear interpolation since $\varepsilon$ is small. The detailed algorithm for computing the approximate Whittle index for our RMAB with general observation models is given below.
@@
9:              for $l=1,...,L$ do Compute $\boldsymbol{\omega^{\prime}}=\mathcal{B}_{l}\boldsymbol{\omega}$
@@
11:                       if $||\omega^{\prime}-\omega^{\prime\prime}||>\varepsilon$ then
@@
Remark. In the adaptive-greedy part of this algorithm, we need to input the parameter $\boldsymbol{A}$ of every arm. For the approximate state space, we can solve for all parameters $A_{\boldsymbol{\omega}}^{\Omega}$ using the dynamic programming equations of $T_{\boldsymbol{\omega}}^{\Omega}$. The output FAIL value is used to judge whether the problem is PCL-indexable w.r.t. $\tilde{\Omega}$. If FAIL=0, the problem is PCL-indexable, otherwise not.

### v4

@@ 2.6 Optimality of Index Policy
In fact, by Theorem 1, if the initial state is $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, then for any $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, $b(\{0\}\cup\Omega)=\frac{1}{1-\beta}-T_{j}^{\Omega^{c}}$. As for $b(\Omega)=\inf\limits_{\pi}\{\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\}$, when we always active the auxiliary arm, then $\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}=0$. Combining this fact with $b(\Omega)\geq 0$, we have $b(\Omega)=0$. Next, we only need to consider whether state 0 is in $\Omega_{*}$ or not:
@@
(ii) If $0\notin\Omega_{*}$, then $\exists N$, when $n>N$, $0\notin\Omega_{2n-1}$. In this case, $b(\Omega_{*})=b(\Omega_{2n-1})=0$. In summary, we obtain $b(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n+1})$. Similar arguments show that $\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n})\leq b(\Omega_{*}^{-})$. ∎
@@
From the weak duality theorem and above discussion, we know that the optimal policy is the priority policy given by $\{\lambda_{\boldsymbol{\omega_{i}}}\}$. It follows from $A_{0}^{\{0\}\cup\Omega}=1$ and $W_{0}^{\{0\}\cup\Omega}=\lambda$ that the priority index for state 0 is $\lambda$. Therefore, for state $\boldsymbol{\omega_{i}}$, when the system’s subsidy $\lambda=\lambda_{\boldsymbol{\omega_{i}}}$, state $\boldsymbol{\omega}$ is precisely at the critical point for taking active or passive action under the optimal policy. Therefore, this two-arm problem is Whittle indexable, and the Whittle index for state $\boldsymbol{\omega_{i}}$ is $\lambda_{\boldsymbol{\omega_{i}}}$. ∎
@@ 3.1 $T$-Step Approximate State Space
If the repeated states are also taken into account, then $|\Omega(T|\boldsymbol{\omega})|=\frac{(H+1)^{T+1}-1}{H}$. In this case, the number of states will increase exponentially with the number of time steps $T$. However, in practical situations, the phenomenon of state repetitions often occurs due to the long time horizon of the underlying stochastic processes in consideration. In other words, along the process of belief state updates, the updated state may have already appeared within a small number of steps from the start or have a short Euclidean distance to a previous state. In such cases, we can assume that the states encountered within a finite number of steps approximate the true belief state space reasonably well. In this sense, the cardinality of the $T$-step state space can be significantly reduced. Formally, we recursively define the $T$-step approximate state space as follows.
@@
For any $\varepsilon>0$, belief state $\boldsymbol{\omega}$ and belief state space $\Omega$, we say $d(\boldsymbol{\omega},\Omega)\geq\varepsilon$, if $\forall\boldsymbol{\omega}^{\prime}\in\Omega$, we have $\boldsymbol{\omega}-\boldsymbol{\omega}^{\prime}||\geq\varepsilon$. Define
$\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})=\begin{cases}\{\mathcal{B}_{h_{1}}\cdots\mathcal{B}_{h_{T}}\boldsymbol{\omega}:0\leq h_{1},\cdots,h_{T}\leq H\text{ and}\\ d(\mathcal{B}_{h_{1}}\cdots\mathcal{B}_{h_{T}}\boldsymbol{\omega},\tilde{\Omega}(T-1,\varepsilon|\boldsymbol{\omega}))\geq\varepsilon\}\cup\tilde{\Omega}(T-1,\varepsilon|\boldsymbol{\omega}),&T\geq 2\\ \{\mathcal{B}_{h}\boldsymbol{\omega}:0\leq h\leq H\text{ and }d(\mathcal{B}_{h}\boldsymbol{\omega},\{\boldsymbol{\omega}\})\geq\varepsilon\}\cup\{\boldsymbol{\omega}\},&T=1\end{cases}.$
@@
Under the definition of $T$-step approximate state space, numerical calculations show that $|\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})|$ is likely to be much smaller than $|\Omega(T|\boldsymbol{\omega})|$, which will greatly reduce the actual number of states that need to be considered for the problem. The following examples illustrate the above concept.
Example 1 (Opportunistic Spectrum Access).
Consider a cognitive radio system with a set of channels and a secondary user. Each channel has 2 states: busy ($S(t)=0$) or idle ($S(t)=1$), which evolves with time as a Markov chain. At each discrete time $t$, the secondary user chooses one channel to sense and observe its state. Observations are error-prone, i.e., a busy channel may be sensed as idle, and vice versa. If the user senses the channel to be idle ($O(t)=1$) and the channel is indeed idle ($S(t)=1$), a unit reward will be accrued. Set the transition matrix of the Markov chain and observation error as follows.
$\mathbf{P}=\begin{pmatrix}0.8&0.2\\ 0.2&0.8\end{pmatrix},\qquad\mathbf{O}=\begin{pmatrix}0.8&0.2\\ 0.2&0.8\end{pmatrix},\qquad\boldsymbol{\omega}_{0}=\boldsymbol{\pi}=(0.5,0.5),$
where $\boldsymbol{\pi}$ is the stationary distribution under $\mathbf{P}$. For $T=6$, the $\varepsilon$-approximate belief space sizes are $|\tilde{\Omega}(6,10^{-2}\mid\boldsymbol{\omega}_{0})|=13$ and $|\tilde{\Omega}(6,10^{-3}\mid\boldsymbol{\omega}_{0})|=14$, whereas $|\Omega(6\mid\boldsymbol{\omega}_{0})|=1{,}093$. For $T=10$, we further compute that $\bigl|\tilde{\Omega}(10,10^{-2}\mid\boldsymbol{\omega}_{0})\bigr|=15$ and $\bigl|\tilde{\Omega}(10,10^{-3}\mid\boldsymbol{\omega}_{0})\bigr|=21$, whereas $\bigl|\Omega(10\mid\boldsymbol{\omega}_{0})\bigr|=88{,}573$.
Example 2 (Opportunistic Spectrum Access - Extended).
Now we consider a 3-state channel model: low SNIR (signal to noise and interference ratio), i.e., $S(t)=0$, middle SNIR ($S(t)=1$), and high SNIR ($S(t)=2$). The user attempts to send more data if the sensed channel is observed to have a higher SNIR. Set the transition matrix of the Markov chain and observation error as follows.
$\mathbf{P}=\begin{pmatrix}0.85&0.10&0.05\\ 0.10&0.80&0.10\\ 0.02&0.08&0.90\end{pmatrix},\qquad\mathbf{O}=\begin{pmatrix}0.88&0.07&0.05\\ 0.15&0.70&0.15\\ 0.04&0.06&0.90\end{pmatrix},\qquad\boldsymbol{\omega}_{0}=\boldsymbol{\pi}=(0.26,0.30,0.44),$
where $\boldsymbol{\pi}$ is the stationary distribution under $\mathbf{P}$. For $T=6$, we obtain $|\tilde{\Omega}(6,10^{-2}\mid\boldsymbol{\omega}_{0})|=24$ and $|\tilde{\Omega}(6,10^{-3}\mid\boldsymbol{\omega}_{0})|=25$, whereas $|\Omega(6\mid\boldsymbol{\omega}_{0})|=5{,}461$. For $T=10$, we obtain $|\tilde{\Omega}(10,10^{-2}\mid\boldsymbol{\omega}_{0})|=36$ and $|\tilde{\Omega}(10,10^{-3}\mid\boldsymbol{\omega}_{0})|=41$, whereas $|\Omega(10\mid\boldsymbol{\omega}_{0})|=1{,}398{,}101$.
Example 3 (Cybersecurity Monitoring and Detection).
Consider a computer network where each time the administrator chooses a section (arm) in the network to monitor and detect the anomaly if any. The arm has 4 states: attacked ($S(t)=0$), malfunctional ($S(t)=1$), overloaded ($S(t)=2$), and normal ($S(t)=3$). Set the transition matrix of the Markov chain and observation error as follows.
$\mathbf{P}=\begin{pmatrix}0.75&0.12&0.08&0.05\\ 0.08&0.72&0.12&0.08\\ 0.05&0.10&0.70&0.15\\ 0.03&0.04&0.08&0.85\end{pmatrix},\qquad\mathbf{O}=\begin{pmatrix}0.88&0.07&0.04&0.01\\ 0.10&0.80&0.07&0.03\\ 0.05&0.10&0.78&0.07\\ 0.02&0.04&0.07&0.87\end{pmatrix},\qquad\boldsymbol{\omega}_{0}=\boldsymbol{\pi}=(0.16,0.21,0.23,0.40),$
where $\boldsymbol{\pi}$ is the stationary distribution under $\mathbf{P}$. For $T=6$, we obtain $\bigl|\tilde{\Omega}(6,10^{-2}\mid\boldsymbol{\omega}_{0})\bigr|=31$ and $\bigl|\tilde{\Omega}(6,10^{-3}\mid\boldsymbol{\omega}_{0})\bigr|=35$, whereas $\bigl|\Omega(6\mid\boldsymbol{\omega}_{0})\bigr|=19{,}531$. For $T=10$, we obtain $\bigl|\tilde{\Omega}(10,10^{-2}\mid\boldsymbol{\omega}_{0})\bigr|=45$ and $\bigl|\tilde{\Omega}(10,10^{-3}\mid\boldsymbol{\omega}_{0})\bigr|=51$, whereas $\bigl|\Omega(10\mid\boldsymbol{\omega}_{0})\bigr|=12{,}207{,}031$.
Lemma 13.
@@
For any given $\varepsilon>0$, $\exists T$, such that $\forall t>T$ and $\boldsymbol{\omega}_{t}\in\Omega(t|\boldsymbol{\omega})$, $\exists\boldsymbol{\omega}_{T}\in\tilde{\Omega}(T|\boldsymbol{\omega})$, such that $\boldsymbol{\omega}_{t}-\boldsymbol{\omega}_{T}||<\varepsilon$, where $||\cdot||$ is the Euclidean norm.
Proof.
Given an arbitrary initial belief state $\boldsymbol{\omega}$, the belief state space grows as a tree in the compact simplex formed by all possible belief states. Therefore, we can cover the simplex by finitely many open balls with diameter equal to $\epsilon$ (Theorem 2.41 in rudin1976principles). For each ball $i$ containing at least one belief state that can be reached by the tree expansion, record the first time $T_{i}$ that the tree reaches a belief state in the ball. The maximum value of all $\{T_{i}\}$ gives the desired $T$. ∎
Lemma 14.
@@
For any given $\varepsilon>0$, $\lim\limits_{T\rightarrow+\infty}|\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})|<+\infty$.
Proof.
By Lemma 13, there exists a $T_{0}$ such that all belief states after that have a distance smaller than $\epsilon$ to $\Omega(T_{0}|\boldsymbol{\omega})$ and $\lim\limits_{T\rightarrow+\infty}|\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})|=|\Omega(T_{0}|\boldsymbol{\omega})|$ which is finite. ∎
By Lemmas 13 and 14, for a sufficiently large number $T$ of decision epochs, the belief states can be well approximated by a finite set of belief states within $\tilde{\Omega}(T|\boldsymbol{\omega})$. In practice, we have a numerical way of choosing the best $\epsilon$ and $T$ inspired by the idea of avoiding overfitting in neural network training. Specifically, we can randomly choose $0<\epsilon<1$ and $T\geq 1$ at the start and conduct the Monte-Carlo Simulation over a period longer than $T$ to observe the performance. In the next round, we reduce $\epsilon$ by half and observe the performance. We repeatedly refine $\epsilon$ until the performance is no longer improving. Then we do the same procedure by doubling $T$ in each round. In sum, we alternatively adjust $\epsilon$ and $T$ until the performance has no significant improvement and obtain the best $\epsilon$ and $T$. The detailed procedure is given in Algorithm 1.
Algorithm 1 Adaptive selections of $\epsilon$ and $T$
1:Input: initial $\epsilon\in(0,1)$, $T\geq 1$, performance tolerance $\delta_{\text{perf}}>0$, simulation horizon $H>T$.
2:Definition: let $J(\epsilon,T)$ denote the Monte–Carlo estimate of long-run performance over horizon $H$ under parameters $(\epsilon,T)$.
3:Compute the baseline performance $J_{\text{best}}\leftarrow J(\epsilon,T)$ $\triangleright$ best performance observed so far
4:repeat
5:  Step 1. Refine $\epsilon$ with fixed $T$
6:  $\epsilon_{\text{cand}}\leftarrow\epsilon/2$ $\triangleright$ candidate $\epsilon$
7:  $J_{\epsilon}\leftarrow J(\epsilon_{\text{cand}},T)$
8:  if $J_{\epsilon}-J_{\text{best}}>\delta_{\text{perf}}$ then
9:   $\epsilon\leftarrow\epsilon_{\text{cand}}$, $J_{\text{best}}\leftarrow J_{\epsilon}$, repeat Step 1.
10:  Step 2. Refine $T$ with fixed $\epsilon$
11:  $T_{\text{cand}}\leftarrow 2T$ $\triangleright$ candidate $T$
12:  $J_{T}\leftarrow J(\epsilon,T_{\text{cand}})$
13:  if $J_{T}-J_{\text{best}}>\delta_{\text{perf}}$ then
14:   $T\leftarrow T_{\text{cand}}$, $J_{\text{best}}\leftarrow J_{T}$, repeat Step 2
15:until Neither Step 1 nor Step 2 updates $(\epsilon,T)$
16:Output: selected $(\epsilon,T)$ as the adaptive choice for this problem instance.
@@ 3.2 AG Algorithm for Approximate State Space
Now we are ready to incorporate the approximate state space $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$ into the AG algorithms introduced in nino2001restless based on Theorem 12 and Lemmas 13 and 14. Specifically, for the state update, if the resulting belief state $\boldsymbol{\omega}$ does not fall within $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$, then we look for the belief state $\boldsymbol{\omega^{\prime}}$ in $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega_{0}})$ that has the closest Euclidean distance to $\boldsymbol{\omega}$, and use the Whittle index $\lambda_{\boldsymbol{\omega^{\prime}}}$ of $\boldsymbol{\omega^{\prime}}$ as the Whittle index of $\boldsymbol{\omega}$ or adopt a linear interpolation since $\varepsilon$ is small. The detailed algorithm for computing the approximate Whittle index for our RMAB with general observation models is given below.
@@
9:     for h=1,,Hh=1,...,H do Compute $\boldsymbol{\omega^{\prime}}=\mathcal{B}_{h}\boldsymbol{\omega}$
@@
11:        if $\omega^{\prime}-\omega^{\prime\prime}||>\varepsilon$ then
@@
Remark. In the adaptive-greedy part of this algorithm, we need to input the parameter $\boldsymbol{A}$ of every arm. For the approximate state space, we can solve for all parameters $A_{\boldsymbol{\omega}}^{\Omega}$ using the dynamic programming equations of $T_{\boldsymbol{\omega}}^{\Omega}$. The output FAIL value is used to judge whether the problem is PCL-indexable w.r.t. $\tilde{\Omega}$. If FAIL=0, the problem is PCL-indexable, otherwise not. If the approximate state space over any finite horizon fails the AG algorithm, then we can conclude that the original RMAB with infinite states does not satisfy PCL-indexability since the state combination subset violates the partial conservation law.